There are 2 types of popcorn and 4 types of hotdogs available at a concession stand in a cinema
How many types of food are available at the stand?

Answer by ikleyn(52777)

.

                2 * 4 = 8,     if you combine them in pairs.

                2 + 4 = 6,     if you take them separately.

                2 * 4 + 2 + 4 = 14,   if you consider them as combined and/or separate.